OpenStudy (anonymous):

what is the answer to the square root of 16/49

jimthompson5910 (jim_thompson5910):

What is the square root of 16?

OpenStudy (kainui):

What is the square root of 49?

OpenStudy (anonymous):

\[\sqrt{\frac{a}{b}} = \frac{\sqrt{a}}{\sqrt{b}}\]

OpenStudy (anonymous):

i think this is the answer,,